OpenStudy (anonymous):

lim h -> 0 (cos( (pi/2) + h) - cos( (pi/2) ) ) / h

OpenStudy (anonymous):

@RadEn

OpenStudy (raden):

looks like it is the defined of derivative of cos(x), right ?

OpenStudy (anonymous):

yeah

OpenStudy (raden):

nah, the derivative of cos(x) = - sin(x) :)

OpenStudy (anonymous):

yep :)

OpenStudy (raden):

so, the answer is -sin(x)

OpenStudy (anonymous):

but the answer at the back was -1

OpenStudy (anonymous):

@RadEn

OpenStudy (raden):

opppss.. yes, it wants find up the value of (cos(x))' for x=0 we have (cos(x))' = -sin(x) now put x=0, it is -sin(0) = 0 hmm... i dont know whcih parts i made mistake

OpenStudy (anonymous):

so u dont know how the answer is -1 then?

OpenStudy (raden):

well, i will try again... miss all above :)

OpenStudy (raden):

use the identity : cosA - cosB = -2sin(((A+B)/2)sin((A-B)/2)

OpenStudy (anonymous):

wait let me interpret what you just typed

OpenStudy (raden):

so, cos( (pi/2) + h)-cos(pi/2) = -2sin((pi/2)+h+pi/2)/2)sin(((pi/2)+h-pi/2)/2) = -2sin(pi+h)/2)sin(h/2) = -2sin(pi/2+h/2)sin(h/2) remember the identity : sin(pi/2+x) = cos(x) therefore, it can be = -2cos(h/2)sin(h/2)

OpenStudy (raden):

that's just for the numerator

OpenStudy (anonymous):

i'll just ask my teacher tomorrow you use to many parenthesis i cant understand :\

OpenStudy (raden):

we will forward its answer ...

OpenStudy (anonymous):

can you help me with the next question?

OpenStudy (raden):

didnt u want finish it ...

OpenStudy (anonymous):

i'll just ask my calc. teacher tomorrow, he did a shorter version of it :)

OpenStudy (anonymous):

can you help me with another question? :)

OpenStudy (raden):

i will try :)